OpenStudy (anonymous):

Three fractions that are equivalent to: 12/36 4/5 and 3/12

OpenStudy (anonymous):

Let's try again....I can't type today...

OpenStudy (anonymous):

12/36=6/18=4/12=3/9=2/6=1/3

OpenStudy (anonymous):

4/5=8/10=12/15=16/20=20/25....

OpenStudy (anonymous):

3/12=1/4=6/24=9/36=12/48.....

OpenStudy (anonymous):

Basically multiple or divide by a number in the numerator and the denominator.

OpenStudy (anonymous):

Must be the same number in both....